# Defining implication relation for classical logic

### v11

@@ 7 Natural deduction
The natural deduction for IRL is given by the set of rules below, which is equivalent to the Hilbert-style system for IRL defined in Section 3. Verification of the equivalence is not difficult, but omitted here.
1. $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$toI CP (Restricted, cf. Remark 7.0.1)
@@ 9. $\phi\wedge(\psi\vee\chi)\Rightarrow(\phi\wedge\psi)\vee(\phi\wedge\chi)$ Distribution
10. $\{\phi\to\psi,\phi\to\neg\psi\}\Rightarrow\neg\phi$$\neg$¬I RAA
11. $\{\neg\phi\to\psi,\neg\phi\to\neg\psi\}\Rightarrow\phi$$\neg$¬E Proof by contradiction
[…]
14. $\{\phi\to\psi,\psi\to\phi\}\Rightarrow(\phi\leftrightarrow\psi)$leftrightarrowI Bi-implication introduction
15. $(\phi\leftrightarrow\psi)\Rightarrow(\phi\to\psi)$, $(\phi\leftrightarrow\psi)\Rightarrow(\psi\to\phi)$leftrightarrowE Bi-implication elimination
Remark 7.0.1 (Restriction on CP).
By the definition of proof in Subsection 3.1, any member of proof sequence must be either a theorem (or an assumed theorem) or a result from the same-layer premise(s) by a rule (or by an assumed rule), not allowing any other cases. Attention should be paid to this restriction, especially in nested conditional proofs.
As an example, the proof of $(\phi\vee\psi)\wedge(\phi\to\chi)\wedge(\psi\to\chi)\to\chi$ in the natural deduction is shown as follows:
1 $(\phi\vee\psi)\wedge(\phi\to\chi)\wedge(\psi\to\chi)$ Premise
@@ 5 $\phi\vee\psi\to\chi$ 3,4: “Two-to-one”
From 1–6 the result is obtained by the CP rule.
@@ 8 Extension to first-order logic
Let $x,y,...$, $s,t,...$, $P,Q,...$ and $\phi,\psi,...$ be FOL variables, terms, predicates and formulas, respectively.
@@
    • Universal generalization: $\phi(y)\to\forall x(\phi(x))$ where $y$ is arbitrary.
    • Universal instantiation: $\forall x(\phi(x))\to\phi\{t/x\}$.
@@
    • Existential instantiation: $\exists x(\phi(x))\to\phi\{c/x\}$ where c is unknown so that it must not occur earlier in the proof and it also must not occur in the conclusion of the proof.
    • Existential generalization (redundant): $\phi\{t/x\}\to\exists x(\phi(x))$.
    • “Many-to-one” (redundant): $\forall x(\phi(x)\to\psi)\to(\exists x(\phi(x))\to\psi))$ where $x$ does not occur free in $\psi$. Its converse can be derived.
As an example of proofs, consider quantifier switching $\exists x\forall y(\phi(x,y))\to\forall y\exists x(\phi(x,y))$ that is equivalent to $\exists x\forall y(\phi(x,y))\wedge\neg\forall y\exists x(\phi(x,y))\to\bot$. The proof of the latter is shown below.
1 $\exists x\forall y(\phi(x,y))\wedge\neg\forall y\exists x(\phi(x,y))$ Premise
2 $\exists x\forall y(\phi(x,y))$ 1: Conjunction elimination
3 $\forall y(\phi(a,y))$ 2: Existential instantiation
4 $\neg\forall y\exists x(\phi(x,y))$ 1: Conjunction elimination
5 $\exists y\forall x(\neg\phi(x,y))$ 4: Duality
6 $\forall x(\neg\phi(x,b))$ 5: Existential instantiation

### v12

@@ 7 Natural deduction
The natural deduction for IRL is given by the set of rules below:
1. $(\phi\vdash\psi)\Rightarrow(\vdash\phi\to\psi)$toI CP (Restricted, cf. Remark 7.0.1 below)
@@ 9. $\phi\wedge(\psi\vee\chi)\Rightarrow(\phi\wedge\psi)\vee(\phi\wedge\chi)$ Distribution
14. $\{\phi\to\psi,\psi\to\phi\}\Rightarrow\phi\leftrightarrow\psi$leftrightarrowI Bi-implication introduction
15. $\phi\leftrightarrow\psi\Rightarrow\phi\to\psi$, $\phi\leftrightarrow\psi\Rightarrow\psi\to\phi$leftrightarrowE Bi-implication elimination
Remark 7.0.1 (Restriction on CP).
By the definition of proof in Subsection 3.1, any member of a proof sequence must be either a theorem (or a theorem indicated in an outer layer), or a result from a subset of the premises and previous members in its own layer by a rule (or by a rule indicated in an outer layer), not allowing any other cases. Attention should be paid to this restriction for nested proofs, especially for nested conditional proofs.
Equivalence to Hilbert system
The natural deduction for IRL is equivalent to its Hilbert-style system defined in Section 3.
The primitive rules listed above of the natural deduction are already proved in its Hilbert system (cf. Section 3).
On the other hand, all axioms and primitive rules of the Hilbert-style IRL can be derived in this natural deduction. Some derivations are shown as examples in the following.
Rule 7.0.1 (Commutativity of conjunction).
$\phi\wedge\psi\Rightarrow\psi\wedge\phi.$
Proof.
1     $\phi\wedge\psi$ Premise
[…]
5 $\phi\wedge\psi\vdash\psi\wedge\phi$ 1–4: Definition of conditional proof
6 $\phi\wedge\psi\Rightarrow\psi\wedge\phi$ 5: Definition of derivable rule
The last two lines will be omitted in the subsequent proofs.
Rule 7.0.2 (Associativity of conjunction).
$(\phi\wedge\psi)\wedge\chi\Rightarrow\phi\wedge(\psi\wedge\chi).$
Proof.
1 $(\phi\wedge\psi)\wedge\chi$ Premise
[…]
With the rules of commutativity and associativity of conjunction, conjunction introduction and elimination can be generalized to $\{\phi_{1},...,\phi_{n}\}\Rightarrow\phi_{1}\wedge\cdots\wedge\phi_{n}$ and $\phi_{1}\wedge\cdots\wedge\phi_{n}\Rightarrow\phi_{i}$ for any $n\geq 2$, respectively.
Meta Rule 7.0.1 (Equivalence of the three forms).
Proof.
This result is dependent of (cf. Subsection 3.3): (a) the definition of proof (conditional proof and unconditional proof, cf. Subsection 3.1), (b) the definition of derivable rule (cf. Subsection 3.1), (c) the rule of conditional proof (it is primitive), (d) modus ponens (it is primitive), and (e) rules of commutativity and associativity of conjunction (they are derived above). ∎
With this result, we need not distinguish between a rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ and a theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$ in proving and using them. Specifically, from a conditional proof $\{\phi_{1},...,\phi_{n}\}\vdash\psi$ or an unconditional proof $\vdash\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$, we obtain both the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ and the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$; we can use the rule $\{\phi_{1},...,\phi_{n}\}\Rightarrow\psi$ as the theorem $\phi_{1}\wedge\cdots\wedge\phi_{n}\to\psi$, and vice versa.
Theorem 7.0.1 (Reflexivity).
$\phi\to\phi.$
Proof.
1 $\phi$ Premise
[…]
Rule 7.0.3 (Transitivity).
$\{\phi\to\psi,\psi\to\chi\}\Rightarrow\phi\to\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\chi$ Premise
3 $(\psi\to\chi)\to(\phi\to\chi)$ 1: Suffixing
4 $\phi\to\chi$ 2,3: MP
[…]
$\{\phi\vee\psi,\phi\to\chi,\psi\to\chi\}\Rightarrow\chi.$
[…]
Proof.
1 $\phi\vee\psi$ Premise
2 $\phi\to\chi$ Premise
3 $\psi\to\chi$ Premise
4 $\phi\vee\psi\to\chi$ 2,3: “Two-to-one”
5 $\chi$ 1,4: MP
Rule 7.0.5 (Distribution of $\wedge$ over $\vee$, converse).
$(\phi\wedge\psi)\vee(\phi\wedge\chi)\Rightarrow\phi\wedge(\psi\vee\chi).$
Proof.
1 $(\phi\wedge\psi)\vee(\phi\wedge\chi)$ Premise
2     $\phi\wedge\psi$ Assumption
[…]
7 $\phi\wedge\psi\to\phi\wedge(\psi\vee\chi)$ 2–6: CP
8 $\phi\wedge\chi\to\phi\wedge(\psi\vee\chi)$ 2–7: Similarly
9 $\phi\wedge(\psi\vee\chi)$ 1,7,8: Proof by exhaustion
Rule 7.0.6 (Distribution of $\vee$ over $\wedge$).
$\phi\vee(\psi\wedge\chi)\Rightarrow(\phi\vee\psi)\wedge(\phi\vee\chi).$
Proof.
1 $\phi\vee(\psi\wedge\chi)$ Premise
2     $\phi$ Assumption
[…]
6 $\phi\to(\phi\vee\psi)\wedge(\phi\vee\chi)$ 2–5: CP
7     $\psi\wedge\chi$ Assumption
[…]
13 $\psi\wedge\chi\to(\phi\vee\psi)\wedge(\phi\vee\chi)$ 7–12: CP
14 $(\phi\vee\psi)\wedge(\phi\vee\chi)$ 1,6,13: Proof by exhaustion
Rule 7.0.7 (Distribution of $\vee$ over $\wedge$, converse).
$(\phi\vee\psi)\wedge(\phi\vee\chi)\Rightarrow\phi\vee(\psi\wedge\chi).$
Proof.
1 $(\phi\vee\psi)\wedge(\phi\vee\chi)$ Premise
2 $((\phi\vee\psi)\wedge\phi)\vee((\phi\vee\psi)\wedge\chi)$ 1: Distribution (of $\wedge$ over $\vee$, primitive)
3     $(\phi\vee\psi)\wedge\phi$ Assumption
[…]
6 $(\phi\vee\psi)\wedge\phi\to\phi\vee(\psi\wedge\chi)$ 3–5: CP
7     $(\phi\vee\psi)\wedge\chi$ Assumption
8     $(\phi\wedge\chi)\vee(\psi\wedge\chi)$ 7: Distribution (of $\wedge$ over $\vee$, primitive)
9       $\phi\wedge\chi$ Assumption
[…]
12     $\phi\wedge\chi\to\phi\vee(\psi\wedge\chi)$ 9–11: CP
13     $\psi\wedge\chi\to\phi\vee(\psi\wedge\chi)$veeI
14     $\phi\vee(\psi\wedge\chi)$ 8,12,13: Proof by exhaustion
15 $(\phi\vee\psi)\wedge\chi\to\phi\vee(\psi\wedge\chi)$ 7–14: CP
16 $\phi\vee(\psi\wedge\chi)$ 2,6,15: Proof by exhaustion
Rule 7.0.8 (Compatibility of $\to$ with $\wedge$).
$\phi\to\psi\Rightarrow\phi\wedge\chi\to\psi\wedge\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\wedge\chi\to\phi$wedgeE
3 $\phi\wedge\chi\to\psi$ 2,1: Transitivity
4 $\phi\wedge\chi\to\chi$wedgeE
5 $\phi\wedge\chi\to\psi\wedge\chi$ 3,4: “One-to-two”
Rule 7.0.9 (Compatibility of $\to$ with $\vee$).
$\phi\to\psi\Rightarrow\phi\vee\chi\to\psi\vee\chi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\psi\to\psi\vee\chi$veeI
3 $\phi\to\psi\vee\chi$ 1,2: Transitivity
4 $\chi\to\psi\vee\chi$veeI
5 $\phi\vee\chi\to\psi\vee\chi$ 3,4: “Two-to-one”
Theorem 7.0.2 ($\top$ is a theorem).
$\top$.⊤ .
Proof.
[…]
2 $\phi\to\phi$ Reflexivity
3 $\top$ 1,2: MP
Theorem 7.0.3 (LEM, theorem form).
$\phi\vee\neg\phi.$
Proof.
[…]
2 $\top$ $\top$ is a theorem
3 $\phi\vee\neg\phi$ 1,2: MP
Theorem 7.0.4 (ECQ, a form without $\bot$).
$\phi\wedge\neg\phi\to\psi.$
Proof.
1 $\phi\wedge\neg\phi$ Premise
[…]
Theorem 7.0.5 (Dual of ECQ, a form without $\top$).
$\phi\to\psi\vee\neg\psi.$
Proof.
1 $\phi$ Premise
[…]
Rule 7.0.10 (DNI).
$\phi\Rightarrow\neg\neg\phi.$
Proof.
1 $\phi$ Premise
2 $\neg\phi\vee\neg\neg\phi$ LEM
[…]
4 $(\phi\wedge\neg\phi)\vee(\phi\wedge\neg\neg\phi)$ 3: Distribution
5 $\phi\wedge\neg\phi\to\neg\neg\phi$ ECQ
6 $\phi\wedge\neg\neg\phi\to\neg\neg\phi$wedgeE
7 $\neg\neg\phi$ 4,5,6: Proof by exhaustion
Rule 7.0.11 (DNE).
$\neg\neg\phi\Rightarrow\phi.$
Proof.
1 $\neg\neg\phi$ Premise
2 $\phi\vee\neg\phi$ LEM
[…]
4 ($(\neg\neg\phi\wedge\phi)\vee(\neg\neg\phi\wedge\neg\phi)$ 3: Distribution
5 $\neg\neg\phi\wedge\phi\to\phi$wedgeE
6 $\neg\neg\phi\wedge\neg\phi\to\phi$ ECQ
7 $\phi$ 4,5,6: Proof by exhaustion
Rule 7.0.12 (Contraposition, negation introduced).
$\phi\to\psi\Rightarrow\neg\psi\to\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\neg\psi\to\neg\psi$ Reflexivity
3 $\neg\psi\to\phi\vee\neg\phi$ Dual of ECQ
4 $\neg\psi\to\neg\psi\wedge(\phi\vee\neg\phi)$ 2,3: “One-to-two”
5 $\phi\vee\neg\phi\to\psi\vee\neg\phi$ 1: Compatibility of $\to$ with $\vee$
6 $\neg\psi\wedge(\phi\vee\neg\phi)\to\neg\psi\wedge(\psi\vee\neg\phi)$ 5: Compatibility of $\to$ with $\wedge$
7 $\neg\psi\wedge(\psi\vee\neg\phi)\to(\neg\psi\wedge\psi)\vee(\neg\psi\wedge\neg\phi)$ Distribution
8 $\neg\psi\wedge\psi\to\neg\psi\wedge\neg\phi$ ECQ
9 $\neg\psi\wedge\neg\phi\to\neg\psi\wedge\neg\phi$ Reflexivity
10 ($(\neg\psi\wedge\psi)\vee(\neg\psi\wedge\neg\phi)\to\neg\psi\wedge\neg\phi$ 8,9: “Two-to-one”
11 $\neg\psi\wedge\neg\phi\to\neg\phi$wedgeE
12 $\neg\psi\to\neg\phi$ 4,6,7,10,11: Transitivity
Rule 7.0.13 (Contraposition, negation eliminated).
$\neg\phi\to\neg\psi\Rightarrow\psi\to\phi.$
Proof.
1 $\neg\phi\to\neg\psi$ Premise
2 $\psi\to\neg\neg\psi$ DNI
3 $\neg\neg\psi\to\neg\neg\phi$ 1: Contraposition, negation introduced
4 $\neg\neg\phi\to\phi$ DNE
5 $\psi\to\phi$ 2,3,4: Transitivity
Rule 7.0.14 (Prefixing).
$\phi\to\psi\Rightarrow(\chi\to\phi)\to(\chi\to\psi).$
Proof.
1 $\phi\to\psi$ Premise
2 $(\chi\to\phi)\to(\neg\phi\to\neg\chi)$ Contraposition
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 ($(\neg\phi\to\neg\chi)\to(\neg\psi\to\neg\chi)$ 3: Suffixing
5 ($(\neg\psi\to\neg\chi)\to(\chi\to\psi)$ Contraposition
6 $(\chi\to\phi)\to(\chi\to\psi)$ 2,4,5: Transitivity
Rule 7.0.15 (RAA, another $\neg$¬I).
$\{\phi\to\psi,\phi\to\neg\psi\}\Rightarrow\neg\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\neg\psi$ Premise
3 $\neg\psi\to\neg\phi$ 1: Contraposition
4 $\neg\neg\psi\to\neg\phi$ 2: Contraposition
5 $\neg\psi\vee\neg\neg\psi\to\neg\phi$ 3,4: “Two-to-one”
6 $\neg\psi\vee\neg\neg\psi$ LEM
7 $\neg\phi$ 5,6: MP
Rule 7.0.16 (Definition of $\to$ by $\leftrightarrow$).
$\phi\to\psi\Leftrightarrow\phi\wedge\psi\leftrightarrow\phi.$
Proof.
1 $\phi\to\psi$ Premise
2 $\phi\to\phi$ Reflexivity
3 $\phi\to\phi\wedge\psi$ 1,2: “One-to-two”
4 $\phi\wedge\psi\to\phi$wedgeE
[…]
1 $\phi\wedge\psi\leftrightarrow\phi$ Premise
[…]
3 $\phi\wedge\psi\to\psi$wedgeE
4 $\phi\to\psi$ 2,3: Transitivity
Based on some of the above results, it is not difficult to prove the primitive rule $\{\phi\leftrightarrow\psi,\chi\}\Rightarrow\chi[\phi\mapsto\psi]$ of replacement of the Hilbert-style IRL (cf. Subsection 3.2) by induction.
@@ 8 Extension to first-order logic
Let $\{x,y,...\}$, $\{s,t,...\}$, $\{P,Q,...\}$ and $\{\phi,\psi,...\}$ be the FOL variables, terms, predicates and formulas, respectively.
@@
    • Universal generalization: $\phi(y)\to\forall x(\phi(x))$ if $y$ is arbitrary.
    • Universal instantiation: $\forall x(\phi(x))\to\phi\{t/x\}$ where $t$ can be any term.
@@
    • Existential instantiation: $\exists x(\phi(x))\to\phi\{t_{0}/x\}$ where $t_{0}$ is unknown, so it must not occur earlier in the proof and it also must not occur in the conclusion of the proof.
    • Existential generalization (redundant): $\phi\{t_{0}/x\}\to\exists x(\phi(x))$ where $t_{0}$ is some (known) term.
    • “Many-to-one” (redundant): $\forall x(\phi(x)\to\psi)\to(\exists x(\phi(x))\to\psi))$ where $x$ does not occur free in $\psi$. Its converse can be derived so is redundant as well.
The following are some derived results.
Theorem 8.0.1 (“One-to-many”, converse).
$(\phi\to\forall x(\psi(x)))\to\forall x(\phi\to\psi(x))$
where $x$ does not occur free in $\phi$.
Proof.
1 $\phi\to\forall x(\psi(x))$ Premise
2 $\forall x(\psi(x))\to\psi(y)$ Universal instantiation ($y$ is arbitrary)
3 $\phi\to\psi(y)$ 1,2: Transitivity
4 $\forall x(\phi\to\psi(x))$ 3: Universal generalization (as $y$ is arbitrary from line 2)
Theorem 8.0.2 (Duality, existential to universal quantifier).
$\neg\exists x(\phi)\to\forall x(\neg\phi).$
Proof.
1 $\neg\exists x(\phi)$ Premise
2 $\neg\forall x(\neg\phi)\to\exists x(\neg\neg\phi)$ Duality, universal to existential quantifier
3 $\neg\forall x(\neg\phi)\to\exists x(\phi)$ 2: Double negation
4 $\neg\neg\forall x(\neg\phi)$ 1,3: Modus tollens
5 $\forall x(\neg\phi)$ 4: Double negation
Theorem 8.0.3 (Existential generalization).
$\phi\{t_{0}/x\}\to\exists x(\phi(x))$
where $t_{0}$ is some (known) term.
Proof.
1 $\phi\{t_{0}/x\}$ Premise
2 $\forall x(\neg\phi(x))\to\neg\phi\{t_{0}/x\}$ Universal instantiation
3 $\neg\forall x(\neg\phi(x))$ 1,2: Modus tollens
4 $\exists x(\neg\neg\phi(x))$ 3: Duality
5 $\exists x(\phi(x))$ 4: Double negation
Theorem 8.0.4 (“Many-to-one”).
$\forall x(\phi(x)\to\psi)\to(\exists x(\phi(x))\to\psi))$
where $x$ does not occur free in $\psi$.
Proof.
1 $\forall x(\phi(x)\to\psi)$ Premise
2 $\phi(y)\to\psi$ 1: Universal instantiation ($y$ is arbitrary)
3 $\neg\psi\to\neg\phi(y)$ 2: Contraposition
4 $\forall x(\neg\psi\to\neg\phi(x))$ 3: Universal generalization (as $y$ is arbitrary from line 2)
5 $\neg\psi\to\forall x(\neg\phi(x))$ 4: “One-to-many”
6 $\neg\forall x(\neg\phi(x))\to\psi$ 5: Contraposition, Double negation
7 $\exists x(\phi(x))\to\psi$ 6: Duality
Theorem 8.0.5 (“Many-to-one”, converse).
$(\exists x(\phi(x))\to\psi)\to\forall x(\phi(x)\to\psi)$
where $x$ does not occur free in $\psi$.
Proof.
1 $\exists x(\phi(x))\to\psi$ Premise
2 $\neg\psi\to\neg\exists x(\phi(x))$ 1: Contraposition
3 $\neg\psi\to\forall x(\neg\phi(x))$ 2: Duality
4 $\forall x(\neg\psi\to\neg\phi(x))$ 3: “One-to-many”, converse
5 $\neg\psi\to\neg\phi(y)$ 4: Universal instantiation ($y$ is arbitrary)
6 $\phi(y)\to\psi$ 5: Contraposition
7 $\forall x(\phi(x)\to\psi)$ 6: Universal generalization (as $y$ is arbitrary)
Theorem 8.0.6 (Universal conjunction).
$\forall x(\phi(x)\wedge\psi(x))\leftrightarrow\forall x(\phi(x)\wedge\forall x(\psi(x).$
Proof.
1 $\forall x(\phi(x)\wedge\psi(x))$ Premise
2 $\phi(y)\wedge\psi(y)$ 1: Universal instantiation ($y$ is arbitrary)
3 $\phi(y)$ 2: Conjunction elimination
4 $\forall x(\phi(x))$ 3: Universal generalization (as $y$ is arbitrary)
5 $\forall x(\psi(x))$ 2–4: Similarly
6 $\forall x(\phi(x)\wedge\forall x(\psi(x)$ 4,5: Conjunction introduction
The converse can be proved in the reverse order. ∎
Theorem 8.0.7 (Existential disjunction).
$\exists x(\phi(x)\vee\psi(x))\leftrightarrow\exists x(\phi(x)\vee\exists x(\psi(x)).$
Proof.
1 $\exists x(\phi(x)\vee\psi(x))$ Premise
2 $\phi(a)\vee\psi(a)$ 1: Existential instantiation
3 $\phi(a)\to\exists x(\phi(x))$ 2: Existential generalization
4 $\phi(a)\to\exists x(\phi(x))\vee\exists x(\psi(x))$ 3: Weakening
5 $\psi(a)\to\exists x(\phi(x))\vee\exists x(\psi(x))$ 2–4: Similarly
6 $\exists x(\phi(x))\vee\exists x(\psi(x))$ 2,4,5: Proof by exhaustion
1 $\exists x(\phi(x))\vee\exists x(\psi(x))$ Premise
2 $\exists x(\phi(x))\to\phi(a)$ Existential instantiation
3 $\exists x(\phi(x))\to\phi(a)\vee\psi(a)$ 2: Weakening
4 $\phi(a)\vee\psi(a)\to\exists x(\phi(x)\vee\psi(x))$ Existential generalization
5 $\exists x(\phi(x))\to\exists x(\phi(x)\vee\psi(x))$ 3–4: Transitivity
6 $\exists x(\psi(x))\to\exists x(\phi(x)\vee\psi(x))$ 2–5: Similarly
7 $\exists x(\phi(x)\vee\psi(x))$ 1,5,6: Proof by exhaustion
Theorem 8.0.8 (Quantifier switching).
$\exists x\forall y(\phi(x,y))\to\forall y\exists x(\phi(x,y)).$
Proof.
1     $\exists x\forall y(\phi(x,y))$ Assumption
2     $\neg\forall y\exists x(\phi(x,y))$ Assumption
3     $\forall y(\phi(a,y))$ 1: Existential instantiation
4     $\exists y\forall x(\neg\phi(x,y))$ 2: Duality
5     $\forall x(\neg\phi(x,b))$ 4: Existential instantiation